# On Effects of Steering Latent Representation for Large Language Model Unlearning

### v1

@@ 4.1 Measuring token confidence with MaxLogit
As discussed in Section 3.1, we test our hypothesis by considering the Maximum Logit Value (MaxLogit) estimator for measuring the token confidence. More specifically, we compute the MaxLogit for each token $x_{n+1}$ given a sequence of tokens $x_{1:n}=\{x_{1},...,x_{n}\}$ from vocabulary $V$ as:
$\text{MaxLogit}(x_{n+1})=\max_{x_{n+1}\in}f^{\text{unlearn}}(x_{n+1}|x_{1:n}),$ (25)
We use WMDP-Biology and WMDP-Cyber Q&A datasets (Li et al. 2024b) with total $3260$ Q&As. We formulated each question and answer as a default zero-shot Q&A prompt to query the unlearned LLM (Gao et al. 2023). The detail of the prompt template are located in Appendix A.1. We used greedy decoding to generate tokens and compute the MaxLogit of each token over $k=30$ generated tokens. The MaxLogit distribution was then analyzed for each model Base vs. RMU (unlearned on WMDP-Biology and WMDP-Cyber forget datasets).
The results are presented in Fig. 2 (a)-(d). We find that the MaxLogit distribution for the base model is generally wider compared to the RMU model. In contrast, the RMU model demonstrates a more concentrated and approximately normal distribution of MaxLogit values. The peak of the RMU model’s MaxLogit distribution is shifted towards lower values relative to the base model. This indicates that the RMU model tends to assign lower confidence scores to the generated tokens. Overall, the RMU model’s MaxLogit distribution exhibits lower compared to the base model, thereby verifying our analysis.
@@ 4.2 The effect of the coefficient $c$
We analyze the impact of $c$ for forgotten knowledge and retained knowledge, using WMDP (Li et al. 2024b) and MMLU (Hendrycks et al. 2020). See Section 6 for the full experiment setting. Fig. 3a shows: (i) a clear positive correlation between the drop-in-accuracy rate and the value of $c$, i.e. higher $c$ makes the accuracy decrease faster. (ii) A larger value of $c$ tends to make a more drop-in-accuracy on WMDP (Fig. 3a). (iii) However, a larger $c$ comes with a caveat in a significant drop in general performance on MMLU (Fig. 3b).
@@
Figure 3: Average accuracy of (a) WMDP (Biology and Cyber) and (b) MMLU-All with different coefficient $c$.
@@ 4.3 The effect of layers on unlearning
We investigate the effect of unlearn layers on accuracy and the representation norm during unlearning. We change the unlearn layer $l$ from $3\to 31$, fixed $c=6.5$. Table 1 shows that RMU is effective for unlearning within the early layers ($3\to 10$), yet exhibits inefficacy within middle and later layers ($11\to 31$). Interestingly, in Fig. 4, we observed that within early layers, the $\ell^{2}$-norm of forget samples are smaller than the coefficient $c$. During unlearning, the representation norm exponentially increases, approaching $c$, thereby facilitating the convergence of forget loss. Conversely, within middle and later layers, the representation norms of forget samples, initially larger than $c$, remain unchanged during unlearning, making the forget loss divergent
@@ 5 Adaptive RMU
Task/unlearn layer base 3 4 5 6 7 8 9 10 11 12 13 14 15 16
WMDP-Biology $\downarrow$ 63.7 31.3 42.2 34.8 29.3 28.8 36.6 41.1 50.9 62.7 59.2 62.1 63.2 63.0 64.1
WMDP-Cyber $\downarrow$ 43.5 43.0 42.1 31.0 27.8 28.8 30.4 29.1 29.8 37.2 39.5 38.4 41.8 42.4 43.4
MMLU $\uparrow$ 58.1 57.2 56.8 57.0 57.0 56.8 56.8 57.2 57.9 57.7 57.3 57.2 57.9 58.3 57.9
Average $\uparrow$ 36.8 34.1 38.8 41.0 40.8 38.4 37.8 35.5 30.6 30.7 30.2 29.5 29.6 28.8
Task/unlearn layer 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31
WMDP-Biology $\downarrow$ 63.7 63.8 63.7 63.2 63.4 63.7 63.6 63.8 63.7 63.5 63.5 63.5 63.7 63.5 64.0
WMDP-Cyber $\downarrow$ 43.5 44.1 43.7 43.8 43.9 43.9 43.7 43.5 43.4 43.8 43.6 43.8 43.7 43.7 43.9
MMLU $\uparrow$ 57.9 58.1 58.1 58.1 58.1 58.0 58.0 58.0 58.1 58.1 58.1 58.0 58.1 58.0 58.0
Average $\uparrow$ 28.9 28.8 29.0 29.1 29.0 28.9 28.9 28.9 29.0 29.0 29.0 28.9 29.0 29.0 28.8
Table 1: Q&A accuracy of RMU Zephyr-7B models on WMDP-Biology, WMDP-Cyber, and MMLU w.r.t unlearn layer $l$ from $3\to 31$. The coefficient $c=6.5$. The best and runner up are marked.
Task/unlearn layer base 3 4 5 6 7 8 9 10 11 12 13 14 15 16
WMDP-Biology $\downarrow$ 63.7 30.9 29.7 25.8 27.1 23.7 24.3 24.6 27.1 38.8 30.2 35.1 51.3 31.7 39.5
WMDP-Cyber $\downarrow$ 43.5 43.2 38.9 24.4 24.3 26.5 25.2 27.0 27.1 27.8 27.0 27.0 27.4 29.3 29.1
MMLU $\uparrow$ 58.1 56.8 56.1 55.0 55.1 55.0 54.0 50.4 55.9 54.0 47.6 40.9 56.7 55.5 57.3
Average $\uparrow$ 36.6 37.7 41.8 41.5 41.7 41.4 39.1 41.2 37.1 36.3 31.7 35.4 39.3 38.3
Task/unlearn layer 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31
WMDP-Biology $\downarrow$ 44.1 37.3 47.6 46.7 49.4 49.6 51.3 55.2 53.0 58.9 47.6 64.1 58.7 56.3 64.8
WMDP-Cyber $\downarrow$ 31.1 26.8 26.6 26.8 27.2 27.8 28.0 36.4 37.8 43.8 43.3 43.9 42.2 43.9 44.0
MMLU $\uparrow$ 57.4 57.4 56.8 56.9 57.8 57.8 57.6 57.9 57.8 57.8 57.6 58.2 57.9 58.0 58.0
Average $\uparrow$ 36.7 39.4 36.6 36.8 36.5 36.3 35.7 32.8 33.0 30.0 32.8 28.9 30.5 30.7 31.0
Table 2: Q&A accuracy of Adaptive RMU Zephyr-7B models on WMDP-Biology, WMDP-Cyber, and MMLU w.r.t unlearn layer $l$ from $3\to 31$. The scaling factor $\beta=5$. The best and runner up are marked.
Inspired by the observations in Section 4.3, we propose Adaptive RMU, a simple yet effective alternative method with an adaptive forget loss by scaling the random unit vector $\bm{u}$ with an adaptive scaling coefficient $\beta||h_{\theta^{\text{frozen}}}^{(l)}(x_{F})||_{2}$, where $\beta\in\mathbb{R}^{+}$ is a scaling factor and $||h_{\theta^{\text{frozen}}}^{(l)}(x_{F})||_{2}$ is the $\ell^{2}$-norm of forget samples $x_{F}$ on model $f_{\theta^{\text{frozen}}}$. The total loss is calculated as follows:
$\mathcal{L}^{\textnormal{adap}}$ $=\underbrace{||h_{\theta^{\text{unlearn}}}^{(l)}(x_{F})-\beta||h_{\theta^{\text{frozen}}}^{(l)}(x_{F})||_{2}\cdot\bm{u}||_{2}^{2}}_{\textnormal{adaptive forget loss}}$
$+\alpha\underbrace{||h_{\theta^{\text{unlearn}}}^{(l)}(x_{R})-h_{\theta^{\text{frozen}}}^{(l)}(x_{R})||_{2}^{2}}_{\textnormal{retain loss}}$ (26)
Our Adaptive RMU is shown in Algorithm 1. In Appendix A.2, we show that Adaptive RMU has the same computational complexity as RMU.
@@
12:     Compute the adaptive loss $\mathcal{L}^{\textnormal{adap}}$ by Eqn. 26.
@@ 6 Experiment
We use WMDP-Biology and WMDP-Cyber forget datasets as $\mathcal{D}_{\text{forget}}$ and Wikitext (Merity et al. 2016) as $\mathcal{D}_{\text{retain}}$ for fine-tuning the LLM. Unlearned models are evaluated on WMDP Q&A datasets and MMLU (Hendrycks et al. 2020). An unlearned model has a higher average of accuracy on MMLU and drop-in-accuracy on WMDP is better. Details of the datasets can be found in the Appendix A.1.
@@
We use the following LLMs: Zephyr-7B (Tunstall et al. 2023), Meta Llama-3-8B (Meta 2024), and Mistral-7B (Jiang et al. 2023).
@@
Models were fine-tuned using AdamW (Loshchilov and Hutter 2019) with learning rate $\eta=5e-5$, batch-size of $4$, max sequence len of $512$ for WMDP-Biology and $768$ for WMDP-Cyber, with $T=500$ gradient update steps. The retain weight $\alpha=1200$. For the baseline RMU, we follow the previous work and let $c=6.5$ (Li et al. 2024b). We grid search for unlearn layer $l$ from the third layer to the last layer. For the Adaptive RMU, we search for the scaling factor $\beta\in\{2,3,5,10\}$. We update three layers parameters $\{l,l-1,l-2\}$ of the model. Two NVIDIA A40s with 90GB RAM were used to run the experiment.
@@
Method/tasks WMDP-Bio $\downarrow$ WMDP-Cyber $\downarrow$ MMLU $\uparrow$ Average $\uparrow$
@@
LLMU 59.5 39.5 44.7 24.4
SCRUB 43.8 39.3 51.2 31.6
SSD 50.2 35.0 40.7 25.8
RMU ($l=7$) 28.8 28.8 56.8 40.8
Adaptive RMU ($l=7$) 23.7 26.5 55.0 41.7
Table 3: Average of drop-in-accuracy on WMDP and accuracy on MMLU. The best and runner $\underline{\textnormal{runner up}}$ are marked.
Table 1 and 2 show that Adaptive RMU with Zephyr-7B models significantly improves RMU, reducing average accuracy by $13.1\%$ on WMDP-Biology and $3.6\%$ on WMDP-Cyber within early layers ($3\to 10$), and by $15.6\%$ on WMDP-Biology and $9.6\%$ on WMDP-Cyber within middle and later layers ($11\to 31$). This corresponds to an overall enhancement of $14.3\%$ and $6.6\%$ in drop-in-accuracy for the WMDP-Biology and WMDP-Cyber, respectively. Table 3 also shows that Adaptive RMU surpasses RMU, LLMU, SCRUB, and SSD by $0.9\%$, $17.3\%$, $10.2\%$, and $15.9\%$ in term of the average of drop-in-accuracy on WMDP and accuracy on MMLU, respectively, establishing a new state-of-the-art performance. See Appendix B for full results on other models and settings.
@@ 7 Conclusion
Appendix A Datasets and algorithm computational perplexity
@@ A.1 Datasets
WMDP (Li et al. 2024b).
The WMDP benchmark, standing for the Weapon of Mass Destruction Proxy, is a corpora consisting of forget sets, retain sets, and Q&A sets. The WMDP Q&A is a dataset of $3,668$ multiple-choice questions about Biosecurity ($1,273$), Cybersecurity ($1,987$), and Chemical security ($408$). The WMDP-Biology forget and retain sets consist of papers from PubMed. The WMDP-Biology forget comprises papers used in generating WMDP-Biology questions, while the retain set samples papers from various categories within general biology. The retain set excludes papers from the forget set and employs keyword exclusion to avoid topics related to Q&A set. The WMDP-Cyber forget and retain sets consist of passages crawled from GitHub with two different sets of keywords. Note that we did not benchmark for WMDP-Chemistry Q&A due to no WMDP-Chemistry forget set being publicly released. This dataset is available at https://github.com/centerforaisafety/wmdp
MMLU (Hendrycks et al. 2020).
The MMLU benchmark, standing for Massive Multitask Language Understanding, a dataset of $15,908$ multiple-choice Q&A covers $57$ subjects across STEM, the humanities, social science, and more. MMLU is designed to measure general knowledge by evaluating models in zero-shot or few-shot settings. This dataset is available at https://huggingface.co/datasets/cais/mmlu
@@
We follow the default zero-show Q&A format to query LLM (Gao et al. 2023). For example:
@@ D. Neither A nor B
The difference between RMU and Adaptive RMU is the calculation of the forget coefficient. Adaptive RMU uses adaptive coefficient $\beta||h_{\theta^{\text{frozen}}}^{(l)}(x_{F})||_{2}$, which can be calculated and cached during the first iteration of the inner for loop in Algorithm 1. Thus, the complexity of Adaptive RMU is equal to that of RMU. Additionally, we report the average unlearning runtime in Table 4.
Mistral-7B Zephyr-7B Meta Llama-3-8B
1225.2 1254.0 1729.8
Table 4: Average unlearning runtime in second (with 2 NVIDIA A40s, batch-size of 4 and 500 steps update)
@@ B.1 Unlearning performance of other models
We report the unlearning performance of Adaptive RMU Llama-3-8B, and Mistral-7B models in Table 5, and 6. We observed a clear trend that the unlearning performance is more effective when using the early layer as the unlearn layer.
@@
Average $\uparrow$ 43.5 42.2 42.0 34.7 43.6 36.8 36.8 36.1 30.9 32.8 31.1 31.3 30.9 31.8 31.3
Table 5: Q&A accuracy of Adaptive RMU Llama-3-8B models on WMDP-Biology, WMDP-Cyber, and MMLU w.r.t unlearn layer $l$ from $3\to 31$. The scaling factor $\beta=5$. The best and runner up are marked.
@@
Average $\uparrow$ 42.4 32.4 39.7 38.5 39.7 37.1 36.8 31.9 31.8 30.8 30.4 29.6 30.5 31.1 29.6
Table 6: Q&A accuracy of Adaptive RMU Mistral-7B models on WMDP-Biology, WMDP-Cyber, and MMLU w.r.t unlearn layer $l$ from $3\to 31$. The scaling factor $\beta=5$. The best and runner up are marked.
B.2 Unlearning performance on MMLU subset unlearning benchmark
We do additional experiments on the MMLU subset unlearning benchmark with three settings:

### v2

@@ 4.1 Measuring token confidence with MaxLogit
As discussed in Section 3.1, we validate our hypothesis by considering the Maximum Logit Value (MaxLogit) estimator for measuring the token confidence. More specifically, we compute the MaxLogit for each token $x_{n+1}$ given a sequence of tokens $x_{1:n}=\{x_{1},...,x_{n}\}$ from vocabulary $V$ as:
$\text{MaxLogit}(x_{n+1})=\max_{x_{n+1}\in V}f^{\text{unlearn}}(x_{n+1}|x_{1:n})$ (23)
We use WMDP-Biology and WMDP-Cyber Q&A datasets (Li et al. 2024b) with total $3260$ Q&As. We formulated each question and answer as a zero-shot Q&A prompt to query the unlearned LLM. The details of the prompt template are located in Appendix A.1. We used greedy decoding to generate tokens and compute the MaxLogit of each token over $k=30$ generated tokens. The MaxLogit distribution was then analyzed for each model Base vs. RMU (unlearned on WMDP-Biology and WMDP-Cyber forget datasets).
The results are presented in Fig. 2 (a)-(d). We find that the MaxLogit distribution for the base model is generally wider compared to the RMU model. In contrast, the RMU model demonstrates a more concentrated and approximately normal distribution of MaxLogit values. The peak of the RMU model’s MaxLogit distribution is shifted towards lower values relative to the base model. This indicates that the RMU model tends to assign lower confidence scores to the generated tokens. Overall, the RMU model’s MaxLogit distribution exhibits lower compared to the base model.
@@ 4.2 The effect of the coefficient $c$
We analyze the impact of $c$ for forgotten knowledge and retained knowledge, using WMDP (Li et al. 2024b) and MMLU (Hendrycks et al. 2021). See Section 6 for the full experiment setting. Fig. 3a shows: (i) a clear positive correlation between the drop-in-accuracy rate and the value of $c$, i.e. higher $c$ makes the accuracy decrease faster. (ii) A larger value of $c$ tends to make a more drop-in-accuracy on WMDP. (iii) However, a larger $c$ comes with a caveat in a significant drop in general performance on MMLU (Fig. 3b).
@@
Figure 3: Average accuracy of (a) WMDP (Biology and Cyber) and (b) MMLU with different coefficient $c$.
@@ 4.3 The effect of layers on unlearning
We investigate the effect of unlearn layers on accuracy and the representation norm during unlearning. Following original work, we change the unlearn layer $l$ from $3\to 31$, fixed $c=6.5$. Fig. 5 shows that RMU is effective for unlearning within the early layers ($3\to 10$), yet exhibits inefficacy within middle and later layers ($11\to 31$). Interestingly, in Fig. 4, we observed that within early layers, the $\ell^{2}$-norm of forget samples are smaller than the coefficient $c$. During unlearning, the representation norm exponentially increases, approaching $c$, thereby facilitating the convergence of forget loss. Conversely, within middle and later layers, the representation norms of forget samples, initially larger than $c$, remain unchanged during unlearning, making the forget loss non-convergence.
@@ 5 Adaptive RMU
12:     Compute the adaptive loss $\mathcal{L}^{\textnormal{adaptive}}$ by Eqn. 24.
@@
Figure 5: Q&A accuracy of RMU and Adaptive RMU Zephyr-7B models on WMDP-Biology, WMDP-Cyber, and MMLU w.r.t unlearn layer $l$ from $3\to 31$.
Inspired by the observations in Section 4.3, we propose Adaptive RMU, a simple yet effective alternative method with an adaptive forget loss by scaling the random unit vector $\bm{u}$ with an adaptive scaling coefficient $\beta||h_{\theta^{\text{frozen}}}^{(l)}(x_{F})||$, where $\beta\in\mathbb{R}$ is a scaling factor and $||h_{\theta^{\text{frozen}}}^{(l)}(x_{F})||$ is the $\ell^{2}$-norm of forget-sample $x_{F}$ on model $f_{\theta^{\text{frozen}}}$. The total loss is calculated as follows:
$\mathcal{L}^{\textnormal{adaptive}}$ $=\underbrace{\mathbb{E}_{x_{F}\in\mathcal{D}_{\text{forget}}}||h_{\theta^{\text{unlearn}}}^{(l)}(x_{F})-\beta||h_{\theta^{\text{frozen}}}^{(l)}(x_{F})||\bm{u}||_{2}^{2}}_{\textnormal{adaptive forget loss}}$
$+\alpha\underbrace{\mathbb{E}_{x_{R}\in\mathcal{D}_{\text{retain}}}||h_{\theta^{\text{unlearn}}}^{(l)}(x_{R})-h_{\theta^{\text{frozen}}}^{(l)}(x_{R})||_{2}^{2}}_{\textnormal{retain loss}}$ (24)
Our Adaptive RMU is shown in Algorithm 1. We note that Adaptive RMU aims to address the challenge of adaptively determining the coefficient $c$ in RMU. We acknowledge that the introduced value $\beta$ is manually tuned via grid search, leaving the challenge to not fully resolved. However, we emphasize that Adaptive RMU offers significant computational advantages over the original RMU. More concretely, in RMU, grid search is conducted over both $c$ and layer $l$ for $l\in[1...L]$, where $L$ is the number of layers. Our analysis suggests that effective unlearning can be achieved when $c$ is higher than the representation norm of forget-samples. Therefore, given a layer $l$, Adaptive RMU only requires tuning $\beta$, which is $L$ times less than that of RMU. This reduction in computational overhead represents a significant improvement when the size of modern deep networks grows.
@@ 6 Experiment
We use WMDP-Biology and WMDP-Cyber forget datasets as $\mathcal{D}_{\text{forget}}$ and Wikitext (Merity et al. 2022) as $\mathcal{D}_{\text{retain}}$ for unlearning the LLM. Unlearned models are evaluated on WMDP Q&A datasets and MMLU (Hendrycks et al. 2021). Details of the datasets can be found in the Appendix A.1.
@@
We use the following LLMs: Zephyr-7B-$\beta$ (Tunstall et al. 2023), Yi-6B (Young et al. 2024), Meta Llama-3-8B (Meta 2024), and Mistral-7B (Jiang et al. 2023).
@@
Models were fine-tuned using AdamW (Loshchilov and Hutter 2019) with learning rate $\eta=5e-5$, batch-size of $4$, max sequence len of $512$ for WMDP-Biology and $768$ for WMDP-Cyber, with $T=500$ gradient update steps. The retain weight $\alpha=1200$. For the baseline RMU, we follow the previous work and let $c=6.5$. We grid search for unlearn layer $l$ from the third to the last layer. For the Adaptive RMU, we grid search for the scaling factor $\beta\in\{2,3,5,10\}$. We report the performances of Adaptive RMU models with $\beta=5$. We update three layers parameters $\{l,l-1,l-2\}$ of the model. Two NVIDIA A40s with 90GB GPU were used to run the experiments. Our code is available at https://github.com/RebelsNLU-jaist/llm-unlearning.
@@
Method/tasks WMDP-Biology $\downarrow$ WMDP-Cyber $\downarrow$ MMLU $\uparrow$
@@
LLMU 59.5 39.5 44.7
SCRUB 43.8 39.3 51.2
SSD 50.2 35.0 40.7
RMU ($l=7$) 28.8 28.8 56.8
Adaptive RMU ($l=7$) 23.7 26.5 55.0
Table 1: Q&A accuracy of Zephyr-7B models on WMDP and MMLU. The best and runner $\underline{\textnormal{runner up}}$ are marked.
Fig. 5 shows that Adaptive RMU significantly improves unlearning performances. Specifically, Adaptive RMU reduces average accuracy by $13.1\%$ on WMDP-Biology and $3.6\%$ on WMDP-Cyber within early layers ($3\to 10$), and by $15.6\%$ on WMDP-Biology and $9.6\%$ on WMDP-Cyber within middle and later layers ($11\to 31$). This corresponds to an overall enhancement of $14.3\%$ and $6.6\%$ in drop-in-accuracy for the WMDP-Biology and WMDP-Cyber, respectively. Table 1 further highlights that Adaptive RMU ($l=7$) outperforms RMU ($l=7$), LLMU, SCRUB, and SSD, establishing a new state-of-the-art performance. We defer the full results on other models and settings in Appendix B.
@@ 7 Conclusion
Appendix A Datasets and Q&A template
@@ A.1 Datasets
WMDP (Li et al. 2024b)
stands for Weapon of Mass Destruction Proxy, is a corpora consisting of forget sets, retain sets, and Q&A sets. The WMDP Q&A is a dataset of $3,668$ multiple-choice questions about Biosecurity ($1,273$), Cybersecurity ($1,987$), and Chemical security ($408$). The WMDP-Biology forget and retain sets consist of papers from PubMed. The WMDP-Biology forget comprises papers used in generating WMDP-Biology questions, while the retain set samples papers from various categories within general biology. The retain set excludes papers from the forget set and employs keyword exclusion to avoid topics related to Q&A set. The WMDP-Cyber forget and retain sets consist of passages crawled from GitHub with two different sets of keywords. Note that we did not benchmark for WMDP-Chemistry Q&A due to no WMDP-Chemistry forget set being publicly released. This dataset is available at https://github.com/centerforaisafety/wmdp.
MMLU (Hendrycks et al. 2021)
stands for Massive Multitask Language Understanding, a dataset of $15,908$ multiple-choice Q&A covers $57$ subjects across STEM, the humanities, social science, and more. MMLU is designed to measure general knowledge by evaluating models in zero-shot or few-shot settings. This dataset is available at https://huggingface.co/datasets/cais/mmlu.
@@
We follow the default zero-shot Q&A format to query LLM (Gao et al. 2023). For example:
@@ D. Neither A nor B
The difference between RMU and Adaptive RMU is the calculation of the forget coefficient. RMU uses a predefined coefficient $c$. Adaptive RMU uses an adaptive coefficient, $\beta||h_{\theta^{\text{frozen}}}^{(l)}(x_{F})||$, which can be calculated and cached during the first iteration of the inner for loop in Algorithm 1. Thus, the complexity of Adaptive RMU is equal to that of RMU.
@@ B.1 Unlearning performance of other models
We report the unlearning performance of Adaptive RMU Yi-6B, Llama-3-8B, and Mistral-7B models in Table 2, Table 3, and Table 4. We observed a clear trend that the unlearning performance is more effective when using the early layer as the unlearn layer.
Task/unlearn layer base 3 4 5 6 7 8 9 10 11 12 13 14 15 16
WMDP-Biology $\downarrow$ 64.8 65.0 49.9 35.2 27.8 26.1 63.3 26.2 27.1 27.4 27.1 26.0 25.4 27.2 34.8
WMDP-Cyber $\downarrow$ 41.1 40.7 40.5 37.7 28.1 25.5 39.3 25.6 23.9 26.1 23.6 24.3 24.2 24.0 25.5
MMLU $\uparrow$ 60.0 60.1 57.7 59.4 51.4 56.5 59.9 56.8 53.7 48.1 49.3 57.0 55.6 47.7 53.3
Task/unlearn layer 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31
WMDP-Biology $\downarrow$ 30.3 32.2 27.1 31.9 41.0 53.4 50.4 53.2 39.2 46.0 39.0 42.5 41.6 40.5 64.8
WMDP-Cyber $\downarrow$ 25.3 24.4 24.3 24.5 26.7 29.8 33.9 36.2 34.3 34.6 31.4 30.4 39.6 40.8 40.6
MMLU $\uparrow$ 45.4 52.1 56.7 58.2 59.3 59.4 59.6 59.7 59.4 59.7 59.4 59.4 59.5 59.7 60.1
Table 2: Q&A accuracy of Adaptive RMU Yi-6B models on WMDP-Biology, WMDP-Cyber, and MMLU.
@@
Table 3: Q&A accuracy of Adaptive RMU Meta Llama-3-8B models on WMDP-Biology, WMDP-Cyber, and MMLU.
@@
Table 4: Q&A accuracy of Adaptive RMU Mistral-7B models on WMDP-Biology, WMDP-Cyber, and MMLU.
B.2 Performances on MMLU subset unlearning benchmark
We did additional experiments on the MMLU subset unlearning benchmark with three settings: